# Stage-wise Dynamics of Classifier-Free Guidance in Diffusion Models

### v1

@@ 5 Conclusion
We present the first systematic analysis of Classifier-Free Guidance (CFG) under multimodal conditionals, characterizing its dynamics in three stages: early Direction Shift, mid Mode Separation, and late Concentration. The theory explains how diversity is lost: early mean-shift bias suppresses weaker modes, while late-stage contraction reduces intra-class variability. This stage-wise view clarifies long-standing empirical phenomena and provides concrete predictions that align with observations on modern diffusion models. Beyond theory, we also show that reducing guidance in early and late stages while emphasizing the mid stage offers a simple way to mitigate diversity loss.
@@ Appendix B Related Work
Bradley & Nakkiran (2024) prove that when both conditional and unconditional distributions are one-dimensional zero-mean Gaussians, the closed-form solution of CFG deviates from the expected gamma-weighted distribution. Xia et al. (2024) further extend this result to high-dimensional isotropic Gaussians with differing conditional and unconditional parameters, deriving a closed-form output that again diverges from the gamma-weighted distribution. Li et al. (2025b) consider the effect of the covariance structure of Gaussian distributions on the CFG sampling process, and point out that CFG enhances generation quality by amplifying class-specific features while suppressing generic ones.
Building on these findings, Wu et al. (2024) show that when the unconditional distribution is a mixture of approximately orthogonal Gaussians and the conditional distribution is Gaussian, CFG sampling theoretically increases classification confidence while simultaneously decreasing the entropy of the output distribution. Jin et al. (2025) further relax these assumptions to more general Gaussian mixture settings, and reveal that CFG induces norm inflation and anomalous diffusion effects, thereby providing a theoretical explanation for the over-saturation phenomena observed in practice. Finally, Li & Jiao (2025) analyze CFG under considerably weaker assumptions without restricting the distributional form, and demonstrate that the improvement in classification confidence holds only in an average sense, rather than uniformly across all initial conditions.
In summary, with the exception of Li & Jiao (2025), existing works primarily focus on settings where the conditional distribution is unimodal. This restriction limits their ability to explain critical phenomena such as the loss of diversity under large guidance weights. Although Li & Jiao (2025) relax the distributional assumptions considerably, their overly general setting similarly fails to capture the underlying mechanism of the phenomenon.
@@ Appendix C Proof of Theorem and Proposition
$\tilde{\pi}_{k}({\bm{x}}_{t})$ $=\frac{\pi_{k}\,\mathcal{N}\!\Big({\bm{x}}_{t};\,(1-t){\bm{\mu}}_{k},\,\big((1-t)^{2}\sigma^{2}+t^{2}\big){\bm{I}}\Big)}{\sum_{j}\pi_{j}\,\mathcal{N}\!\Big({\bm{x}}_{t};\,(1-t){\bm{\mu}}_{j},\,\big((1-t)^{2}\sigma^{2}+t^{2}\big){\bm{I}}\Big)}.$ (11)
@@
Assume the class-conditional prior is a $K$-component Gaussian mixture with shared covariance $\sigma^{2}{\bm{I}}$ and weights $\{\pi_{k}\}_{k=1}^{K}$, and let $\bar{\bm{\mu}}=\sum_{k}\pi_{k}{\bm{\mu}}_{k}$. Let the reverse probability flow ODE and its CFG-driven variant be driven by the estimators from Lemma 1:
@@
with a shared random initialization ${\bm{x}}_{1}^{(y)}={\bm{x}}_{1}^{((\textsc{CFG}))}\stackrel{{\scriptstyle d}}{{=}}\mathcal{N}(\mathbf{0},{\bm{I}})$ independent of $y$. Then for any $\omega>1$ there exists $t_{e1}\in(0,1)$ such that for all $t\in[t_{e1},1)$,
@@
$p_{0}({\bm{x}}\mid y)=\pi_{1}\,\mathcal{N}({\bm{x}};{\bm{\mu}}_{1},\sigma^{2}{\bm{I}})+\pi_{2}\,\mathcal{N}({\bm{x}};{\bm{\mu}}_{2},\sigma^{2}{\bm{I}}),\quad\|{\bm{\mu}}_{1}\|=\|{\bm{\mu}}_{2}\|,\;\pi_{1}<\pi_{2}.$
@@
$\pi_{1}\,\mathcal{N}\!\big({\bm{x}}_{t};(1-t){\bm{\mu}}_{1},\,(t^{2}+(1-t)^{2}\sigma^{2}){\bm{I}}\big)>\pi_{2}\,\mathcal{N}\!\big({\bm{x}}_{t};(1-t){\bm{\mu}}_{2},\,(t^{2}+(1-t)^{2}\sigma^{2}){\bm{I}}\big)\quad\forall\,t\in[0,t_{s_{2}}].$
@@
where $\bar{\bm{\mu}}_{t}({\bm{x}})=r_{1}(t,{\bm{x}}){\bm{\mu}}_{1}+r_{2}(t,{\bm{x}}){\bm{\mu}}_{2}$ and $r_{k}(t,{\bm{x}})\propto\pi_{k}\,\mathcal{N}({\bm{x}};(1-t){\bm{\mu}}_{k},\sigma_{t}^{2}{\bm{I}}),\ r_{1}+r_{2}=1.$ Setting ${\bm{x}}_{t}=(1-t){\bm{z}}_{t}$ and simplifying gives
@@
$\pi_{1}\,\mathcal{N}\!\bigl({\bm{x}}_{t};(1-t){\bm{\mu}}_{1},\sigma_{t}^{2}{\bm{I}}\bigr)>\pi_{2}\,\mathcal{N}\!\bigl({\bm{x}}_{t};(1-t){\bm{\mu}}_{2},\sigma_{t}^{2}{\bm{I}}\bigr),$
@@
$\pi_{1}\,\mathcal{N}\!\big({\bm{x}}_{t};(1-t){\bm{\mu}}_{1},\,(t^{2}+(1-t)^{2}\sigma^{2}){\bm{I}}\big)<\pi_{2}\,\mathcal{N}\!\big({\bm{x}}_{t};(1-t){\bm{\mu}}_{2},\,(t^{2}+(1-t)^{2}\sigma^{2}){\bm{I}}\big).$
@@
(A) Explicit posterior differences. For a single Gaussian $\mathcal{N}(\mu_{k},\sigma^{2}{\bm{I}})$ we have the closed form
@@
and for the unconditional prior $\mathcal{N}(0,{\bm{I}})$, $\hat{\bm{x}}_{0|t}({\bm{x}})=\gamma_{t}{\bm{x}}$ with $\gamma_{t}=\frac{1-t}{t^{2}+(1-t)^{2}}$. In the mixture model,
@@
$\dot{\Delta}_{t}^{(y)}=-\Big(\tfrac{\alpha_{t}-1}{t}\,{\bm{I}}+\tfrac{\beta_{t}}{t}\,{\bm{G}}_{t}\Big)\Delta_{t}^{(y)},\qquad\dot{\Delta}_{t}^{(\textsc{CFG})}=-\Big(\tfrac{(1-\omega)\gamma_{t}+\omega\alpha_{t}-1}{t}\,{\bm{I}}+\omega\tfrac{\beta_{t}}{t}\,{\bm{G}}_{t}\Big)\Delta_{t}^{(\textsc{CFG})}.$ (19)
@@
We compared three settings: (i) vanilla-CFG with constant guidance $\omega$; (ii) time-varying CFG (TV-CFG), which weakens guidance in early/late steps and strengthens it mid-way; and (iii) interval-CFG (Kynkäänniemi et al., 2024), which executes CFG only within a specific interval. For TV-CFG, let $N$ be the number of NFEs, $\{t_{n}\}_{n=0}^{N-1}$ the timesteps, and $\{\omega_{t_{n}}\}$ the scales. The schedule is
@@
analogous to $\int_{0}^{1}\omega_{t}\,{\rm d}t=\omega$. For interval-CFG, the values of guidance scale within the CFG interval are also determined by normalization. Figure 5 visualizes the schedules.
Figure 5: Vanilla-CFG, interval-CFG and TV-CFG guidance scale settings at $\omega=9$.
Appendix E additional visulization results

### v2

@@ 5 Conclusion
We present the first systematic analysis of Classifier-Free Guidance (CFG) under multimodal conditional distribution assumption, characterizing its dynamics in three stages: early Direction Shift, mid Mode Separation, and late Concentration. The theory explains how diversity is lost: early mean-shift bias suppresses weaker modes, while late-stage contraction reduces intra-class variability. This stage-wise view clarifies long-standing empirical phenomena and provides concrete predictions that align with observations on modern diffusion models. Beyond theory, we also show that reducing guidance in early and late stages while emphasizing the mid stage offers a simple way to mitigate diversity loss.
6 Acknowledgements
The authors are with the Department of Electronic Engineering, Beijing National Research Center for Information Science and Technology, Tsinghua University, Beijing 100084, China. This work was supported by the National Key Research and Development Program of China (Grant No. 2025YFF0515601) and the National Natural Science Foundation of China (NSAF U2230201).
@@ Appendix B Related Work
Bradley and Nakkiran (2024) prove that when both conditional and unconditional distributions are one-dimensional zero-mean Gaussians, the closed-form solution of CFG deviates from the expected gamma-weighted distribution. Xia et al. (2024) further extend this result to high-dimensional isotropic Gaussians with differing conditional and unconditional parameters, deriving a closed-form output that again diverges from the gamma-weighted distribution. Li et al. (2025b) consider the effect of the covariance structure of Gaussian distributions on the CFG sampling process, and point out that CFG enhances generation quality by amplifying class-specific features while suppressing generic ones.
Building on these findings, Wu et al. (2024) show that when the unconditional distribution is a mixture of approximately orthogonal Gaussians and the conditional distribution is Gaussian, CFG sampling theoretically increases classification confidence while simultaneously decreasing the entropy of the output distribution. Jin et al. (2025) further relax these assumptions to more general Gaussian mixture settings, and reveal that CFG induces norm inflation and anomalous diffusion effects, thereby providing a theoretical explanation for the over-saturation phenomena observed in practice. Finally, Li and Jiao (2025) analyze CFG under considerably weaker assumptions without restricting the distributional form, and demonstrate that the improvement in classification confidence holds only in an average sense, rather than uniformly across all initial conditions.
In summary, with the exception of Li and Jiao (2025), existing works primarily focus on settings where the conditional distribution is unimodal. This restriction limits their ability to explain critical phenomena such as the loss of diversity under large guidance weights. Although Li and Jiao (2025) relax the distributional assumptions considerably, their overly general setting similarly fails to capture the underlying mechanism of the phenomenon.
@@ Appendix C Proof of Theorem and Proposition
$\tilde{\pi}_{k}({\bm{x}}_{t})$ $=\frac{\pi_{k}\,\mathcal{N}\!\Big({\bm{x}}_{t};\,(1-t){\bm{\mu}}_{k},\,\big((1-t)^{2}\sigma^{2}+t^{2}\big){\mathbf{I}}\Big)}{\sum_{j}\pi_{j}\,\mathcal{N}\!\Big({\bm{x}}_{t};\,(1-t){\bm{\mu}}_{j},\,\big((1-t)^{2}\sigma^{2}+t^{2}\big){\mathbf{I}}\Big)}.$ (11)
@@
Assume the class-conditional prior is a $K$-component Gaussian mixture with shared covariance $\sigma^{2}{\mathbf{I}}$ and weights $\{\pi_{k}\}_{k=1}^{K}$, and let $\bar{\bm{\mu}}=\sum_{k}\pi_{k}{\bm{\mu}}_{k}$. Let the reverse probability flow ODE and its CFG-driven variant be driven by the estimators from Lemma 1:
@@
with a shared random initialization ${\bm{x}}_{1}^{(y)}={\bm{x}}_{1}^{((\textsc{CFG}))}\stackrel{{\scriptstyle d}}{{=}}\mathcal{N}(\mathbf{0},{\mathbf{I}})$ independent of $y$. Then for any $\omega>1$ there exists $t_{e1}\in(0,1)$ such that for all $t\in[t_{e1},1)$,
@@
$p_{0}({\bm{x}}\mid y)=\pi_{1}\,\mathcal{N}({\bm{x}};{\bm{\mu}}_{1},\sigma^{2}{\mathbf{I}})+\pi_{2}\,\mathcal{N}({\bm{x}};{\bm{\mu}}_{2},\sigma^{2}{\mathbf{I}}),\quad\|{\bm{\mu}}_{1}\|=\|{\bm{\mu}}_{2}\|,\;\pi_{1}<\pi_{2}.$
@@
$\pi_{1}\,\mathcal{N}\!\big({\bm{x}}_{t};(1-t){\bm{\mu}}_{1},\,(t^{2}+(1-t)^{2}\sigma^{2}){\mathbf{I}}\big)>\pi_{2}\,\mathcal{N}\!\big({\bm{x}}_{t};(1-t){\bm{\mu}}_{2},\,(t^{2}+(1-t)^{2}\sigma^{2}){\mathbf{I}}\big)\quad\forall\,t\in[0,t_{s_{2}}].$
@@
where $\bar{\bm{\mu}}_{t}({\bm{x}})=r_{1}(t,{\bm{x}}){\bm{\mu}}_{1}+r_{2}(t,{\bm{x}}){\bm{\mu}}_{2}$ and $r_{k}(t,{\bm{x}})\propto\pi_{k}\,\mathcal{N}({\bm{x}};(1-t){\bm{\mu}}_{k},\sigma_{t}^{2}{\mathbf{I}}),\ r_{1}+r_{2}=1.$ Setting ${\bm{x}}_{t}=(1-t){\bm{z}}_{t}$ and simplifying gives
@@
$\pi_{1}\,\mathcal{N}\!\bigl({\bm{x}}_{t};(1-t){\bm{\mu}}_{1},\sigma_{t}^{2}{\mathbf{I}}\bigr)>\pi_{2}\,\mathcal{N}\!\bigl({\bm{x}}_{t};(1-t){\bm{\mu}}_{2},\sigma_{t}^{2}{\mathbf{I}}\bigr),$
@@
$\pi_{1}\,\mathcal{N}\!\big({\bm{x}}_{t};(1-t){\bm{\mu}}_{1},\,(t^{2}+(1-t)^{2}\sigma^{2}){\mathbf{I}}\big)<\pi_{2}\,\mathcal{N}\!\big({\bm{x}}_{t};(1-t){\bm{\mu}}_{2},\,(t^{2}+(1-t)^{2}\sigma^{2}){\mathbf{I}}\big).$
@@
(A) Explicit posterior differences. For a single Gaussian $\mathcal{N}(\mu_{k},\sigma^{2}{\mathbf{I}})$ we have the closed form
@@
and for the unconditional prior $\mathcal{N}(0,{\mathbf{I}})$, $\hat{\bm{x}}_{0|t}({\bm{x}})=\gamma_{t}{\bm{x}}$ with $\gamma_{t}=\frac{1-t}{t^{2}+(1-t)^{2}}$. In the mixture model,
@@
$\dot{\Delta}_{t}^{(y)}=-\Big(\tfrac{\alpha_{t}-1}{t}\,{\mathbf{I}}+\tfrac{\beta_{t}}{t}\,{\bm{G}}_{t}\Big)\Delta_{t}^{(y)},\qquad\dot{\Delta}_{t}^{(\textsc{CFG})}=-\Big(\tfrac{(1-\omega)\gamma_{t}+\omega\alpha_{t}-1}{t}\,{\mathbf{I}}+\omega\tfrac{\beta_{t}}{t}\,{\bm{G}}_{t}\Big)\Delta_{t}^{(\textsc{CFG})}.$ (19)
@@
We compared four settings: (i) vanilla-CFG with constant guidance $\omega$; (ii) time-varying CFG (TV-CFG), which weakens guidance in early/late steps and strengthens it mid-way; (iii) $\beta$-CFG (Malarz et al., 2025), which utilizes the probability density function (PDF) of a Beta distribution as the weight schedule; and (iv) interval-CFG (Kynkäänniemi et al., 2024), which executes CFG only within a specific interval.
For TV-CFG, let $N$ be the number of NFEs, $\{t_{n}\}_{n=0}^{N-1}$ the timesteps, and $\{\omega_{t_{n}}\}$ the scales. The schedule is
@@
analogous to $\int_{0}^{1}\omega_{t}\,{\rm d}t=\omega$. For interval-CFG and $\beta$-CFG, the values of the guidance scale within the CFG interval are also determined by normalization, and the relevant hyperparameters are set according to the recommendations in the original paper. Figure 5 visualizes the schedules.
Figure 5: Vanilla-CFG, interval-CFG, $\beta$-CFG and TV-CFG guidance scale settings at $\omega=9$.
Diversity is evaluated by sampling 1,000 prompts from the COCO dataset. For each prompt, 16 images are generated using different random seeds. LPIPS features are extracted for all generated images, and diversity is quantified as the mean squared pairwise distance between feature representations. Larger values correspond to higher sample diversity.
Appendix E Ablation Study on Peak Timing
In this section, we investigate the sensitivity of our method to the peak location of the guidance schedule. Our theoretical framework describes a three-stage dynamic: early direction shift, intermediate mode separation, and late contraction. The symmetric TV-CFG schedule used in our main experiments serves as a minimal, theory-aligned instantiation. To demonstrate that the effectiveness relies on the general stage-wise mechanism rather than fine-grained hyperparameter tuning, we conduct an ablation study varying the peak timing.
We adjust the peak position of the schedule to occur at 20%, 40%, 60%, and 80% of the total sampling steps, fixing the total steps at $\text{NFE}=10$ and the guidance scale at $\omega=7$. We evaluate the performance across CLIP, IR, FID, saturation, and Diversity metrics.
Table 3: Ablation study on the sensitivity of the peak timing location. Experiments were conducted with $\text{NFE}=10$ and $\omega=7$. The results demonstrate robustness when the peak lies within the intermediate regime.
Peak Timings CLIP $\uparrow$ IR $\uparrow$ FID $\downarrow$ Saturation Diversity $\uparrow$
20% 0.317 0.752 31.669 0.401 1.2214
40% 0.319 0.940 28.770 0.279 1.2309
60% 0.319 0.935 27.722 0.229 1.1963
80% 0.319 0.896 27.092 0.201 1.1622
The quantitative results are presented in Table 3. We observe the following:
  • Stability in the Intermediate Regime: When the peak is located within the intermediate stage (40%–60%), the performance remains stable and optimal across most metrics. Notably, CLIP scores saturate at their highest value (0.319), and IR reaches its peak at 40%.
  • Impact of Extreme Timings: Setting the peak too early (20%) or too late (80%) leads to expected degradation in specific areas. An early peak (20%) results in lower IR and higher FID, suggesting that applying strong guidance too soon hinders the quality of the samples. Conversely, a late peak (80%) yields the lowest Diversity (1.1622) and Saturation, consistent with the theoretical prediction that delayed guidance restricts the available generation space.
These findings confirm that the primary factor for success is the “low $\to$ high $\to$ low” structural shape predicted by our three-stage theory. The method is robust to the exact peak location, provided it resides within the mode-separation stage. This supports our conclusion that the stage-wise mechanism, rather than precise hyperparameter selection, governs the behavior of guided sampling.
Appendix F additional visulization results